# Magnetic Skyrmions: from lumps to supercompactons

### v2

@@ Abstract
The magnetic Skyrmion is described by one control parameter and one length scale. We study the two extreme limits of the control parameter – infinitely large and vanishing – and find that the magnetic Skyrmion becomes a “restricted” magnetic Skyrmion and an $\operatorname{O}(3)$ sigma model lump, respectively. Depending on the potential under consideration, the restricted limit manifests differently. In the case of the Zeeman term, the restricted magnetic Skyrmion becomes a “supercompacton” that develops a discontinuity, whereas for the Zeeman term to the power 3/2 it becomes a normal compacton. In both the lump and the restricted limit the solution is given in exact explicit form. We observe that the case of the Zeeman term squared, which can also be understood as a special combination of the Zeeman term and the easy-plane potential – realizable in the laboratory, the analytically exact solution for all values of the coupling – including the BPS case – is also of the lump type. Finally, we notice that certain materials (e.g. Fe1–xCoxSi or Mn1–xFexGe) have a rather large control parameter $\epsilon$ of order 100, making the restricted limit a suitable rough approximation.
@@ 1 Introduction
In this work we consider the theory of magnetic Skyrmions in two spatial dimensions (2D) and apply certain techniques used in the past in the context of the baby-Skyrme model [42]. We show that a dimensionless parameter enables us to control the relation between the Dirichlet term, the one with two derivatives, and the rest of the energy functional that contains the potential and the DM term. By changing this parameter we can flow from an almost BPS theory, close to the $\operatorname{O}(3)$ sigma model which admits lumps as solutions, to a “restricted” magnetic Skyrmion model with just $E_{1}$ and $E_{0}$. The restricted model is analytically solvable. In the case of the ordinary Zeeman potential, which is quadratic in the perturbative excitation fields near the vacuum, a feature never encountered before in the context of the baby-Skyrme model is present: the solution is not only a compacton111 Compactons are solitonic solutions with compact support: they have a nontrivial profile function for $r<R$ and vanishing energy outside, i.e. for $r>R$. Often the fields are continuous but their derivatives are not at $r=R$. Compactons exist in the baby-Skyrme model [43, 44, 45], the BPS-Skyrme model [46], as Q-balls in the gauged (ungauged) $\mathbb{C}P^{N}$ model [47]([48]) and in K-field theories [49], etc. but it is also discontinuous. We may call it a “supercompacton”. The discontinuity is resolved as soon as we turn on a small coefficient for the Dirichlet term. We also study another subtlety of the model; that is, the DM term can be integrated by parts without changing the equation of motion or the energy when the potential is the Zeeman term (giving rise to exponential falloff). However, it turns out that for potentials that give rise to massless perturbative excitations, e.g. the Zeeman squared term, which can be understood as a specific linear combination of the Zeeman term and the easy-plane potential, discarding a boundary term and using a different DM term in the energy functional, is crucial for computing the correct sizes of the solutions in the lump limit, i.e. the limit of the control parameter going to zero.
@@
Finally, we observe that this lump-type solution corresponds to the solution of the critically coupled version of magnetic Skyrmions with the Zeeman term squared, as described by Barton-Singer et al[19]. It is not limited to this scenario: it is also the solution to the full second-order variational equation of motion for all values of the coupling [19]. We find that for this potential, discarding the boundary term changes the energy functional to become dependent on $\epsilon$, hence revealing a phase transition between the homogeneous phase and the Skyrmionic phase.
@@
On the experimental side, we also confront the dimensionless parameter with the various experimental realizations of the magnetic Skyrmion and find that the (near) “restricted” magnetic Skyrmion – corresponding to the case of a large control parameter – is actually realized in some materials, like Fe1–xCoxSi or Mn1–xFexGe. We also confirm that it is possible to experimentally realize the Zeeman squared term, by using magnets with a certain anisotropy term and adjusting the magnetic field of the Zeeman term, so as to complete the square. At this critically tuned value of the magnetic field, the exact analytic lump-type solution is thus realizable in the laboratory.
@@
This paper is organized as follows. In Sec. 2 we review the magnetic Skyrmion in the continuum formulation with various potentials in our notation. In Sec. 2.1, we review the hedgehog Ansatz and discuss the symmetry breaking properties leading to the fixing of the phase parameter. In Sec. 2.2, we discuss integration-by-parts identities that do not alter the equations of motion or the energy for the case of the Zeeman term. In Secs. 2.3 and 2.4 we introduce the sigma model lump and restricted limits of the magnetic Skyrmion. In Sec. 2.5 we review the case of the critical coupling and in Sec. 2.6 we show that the lump solution is a solution to the model with the Zeeman term squared for any coupling. In Sec. 2.7 we provide numerical solutions that interpolate between the lump and restricted limits, verifying the analytical results obtained. In Sec. 2.8 we compare the control parameter to experimental parameters of various chiral magnets. We conclude in Sec. 3 with a discussion.
@@
$\mathcal{H}_{\mathrm{H}}=-\sum_{i<j}^{N}\mathbb{J}_{ij}\ \bm{S}_{i}\cdot\bm{S}_{j}\ ,$ (2.1)
where the interaction is only between nearest neighbor’s spins, and $\mathbb{J}_{ij}$ is the exchange interaction constant (which is negative in ferromagnetic systems). We suppose that $\mathbb{J}_{ij}=J$ for all $\{j=i+1\;|\;i,j\in[1,N]\}$ with $N$ a large integer number. In the mean field approximation, a proper dynamical variable is the expectation value of the spins, i.e. the unit magnetization vector $\bm{n}\in S^{2}$. Its dynamics is governed by the continuum Ginzburg-Landau (GL) effective theory (see e.g. Ref. [6]). This is the static energy of the $\operatorname{O}(3)$ sigma model. Furthermore, the Hamiltonian of the Heisenberg model can be supplemented by the potential-like term $-\bm{B}\cdot\bm{n}$ that describes the Zeeman interaction with an external magnetic field, which we choose to be aligned along the third Cartesian axis, $\bm{B}=B\bm{e}_{3}$.
@@
$\mathcal{H}_{\mathrm{DM}}=-\sum_{ij}^{N}\bm{D}_{ij}\cdot(\bm{S}_{i}\times\bm{S}_{j})\ ,$ (2.2)
where $\bm{D}_{ij}$ is the chiral coupling parameter (supposed to be equal to $D$ for all $\{j=i+1\;|\;i,j\in[1,N]$) which depends on the material. The corresponding energy in the continuum limit (see App. A), at the leading order in the long-wavelength approximation, reduces to the functional:
@@
$E[\bm{n}]=\int\mathop{}\!\mathrm{d}^{2}x\left(\frac{1}{2}\partial_{i}\bm{n}\cdot\partial_{i}\bm{n}+\kappa\epsilon_{iab}\partial^{i}n^{a}n^{b}+V_{p}(\bm{n})+\frac{\lambda}{2}(\bm{n}\cdot\bm{n}-1)\right),$ (2.4)
@@
We will later be using the fact that the DM term can be integrated by parts
$E[\bm{n}]=\int\mathop{}\!\mathrm{d}^{2}x\left(\frac{1}{2}\partial_{i}\bm{n}\cdot\partial_{i}\bm{n}+\kappa\sum_{b=1}^{2}\epsilon_{iab}\partial^{i}n^{a}n^{b}-\kappa\epsilon_{ia3}n^{a}\partial^{i}n^{3}+V_{p}(\bm{n})+{\rm t.d.}\right),$ (2.7)
where the total derivative reads ${\rm t.d.}=\kappa\,\partial_{i}(\epsilon_{ia3}n^{a}n^{3})$.
@@
So the meaning of $\bm{T}$ is a pair of orthogonal unit vectors in $\mathbb{R}^{3}$. The functional (2.3) corresponds to the choice
$\bm{T}=\begin{pmatrix}1&0\\ 0&1\\ 0&0\end{pmatrix}\ ,\qquad\bm{N}=\begin{pmatrix}0\\ 0\\ 1\end{pmatrix}\ .$ (2.10)
The equations of motion are
$\partial_{i}^{2}n^{a}+hpN^{a}\left(1-\bm{n}\cdot\bm{N}\right)^{p-1}-2\kappa\epsilon_{aib}\partial^{i}n^{b}-\lambda n^{a}=0\ ,\qquad\quad\bm{n}\cdot\bm{n}=1\ ,$ (2.11)
where the Lagrange multiplier is
@@
$E[\vartheta,\varphi]$ $=\int\mathop{}\!\mathrm{d}^{2}x\,\bigg{(}\frac{1}{2}(\partial_{i}\vartheta)^{2}+\frac{1}{2}\sin^{2}\vartheta(\partial_{i}\varphi)^{2}+h(1-\cos{\vartheta})^{p}$
@@
where t.d. is a total derivative
${\rm t.d.}=\frac{\kappa}{2}\left[\partial_{1}(\sin(2\vartheta)\sin\varphi)-\partial_{2}(\sin(2\vartheta)\cos\varphi)\right].$ (2.17)
@@
where $\ell$ controls the size of the Skyrmion (2.19) while $\epsilon$, the dimensionless parameter of the theory, controls the relation between $E_{2}$ and $E_{1}+E_{0}$. In order to see the scaling behavior of the theory, we can write
@@
$\bm{n}(x_{1},x_{2})\mapsto R_{12}(\zeta)\bm{n}(\cos\zeta x_{1}-\sin\zeta x_{2},\sin\zeta x_{1}+\cos\zeta x_{2})\ ,\qquad\zeta\in[0,2\pi)\ .$ (2.26)
Moreover, there are the two internal parities as a discrete group $P_{\rm spt}\times P_{\rm int}$. The DM interaction is responsible for the breaking to the diagonal group $P_{\rm diag}$. Up to an additional internal rotation of $\pi$, the parity transformations act as:
@@
$G={P}_{\rm diag}\times\operatorname{SO}(2)_{\rm diag}\times T_{2}\ ,$ (2.28)
where $T_{2}$ corresponds to the 2-dimensional translations.
We can now briefly consider the perturbative spectrum of the theory near the vacuum of the potential. In order to do so, we expand the fields as: $\bm{n}\to\bm{N}+\bm{u}$ and insert it into the static energy density (2.3) for the class of potentials (2.5):
$\begin{array}[]{rlc}\mathcal{E}\simeq\frac{1}{2}\partial_{i}\bm{u}\cdot\partial_{i}\bm{u}+V_{p}(\bm{u})+{\rm t.d.}\ ,\end{array}$ (2.29)
where $i,j=1,2$, $E=\int\mathcal{E}\mathop{}\!\mathrm{d}^{2}x$ and t.d. stands for total derivative. The Lagrangian density of the system is:
$\mathcal{L}[\bm{u}]=\dfrac{1}{2}\partial_{0}\bm{u}\cdot\partial_{0}\bm{u}-\mathcal{E}\ ,$ (2.30)
where Lorentz invariance is explicitly broken by the DM term. At the quadratic level the DM term contributes only with a linear term which is also a total derivative, thus the spectrum of the perturbative particles is not changed from the $\operatorname{O}(3)$ sigma model with the potential $V_{p}$ [50]. When $\bm{T}$ and $\bm{N}$ are not orthogonal there may be a quadratic contribution from the DM term.
@@ 2.1 Hedgehog Solution
[…]
@@
$Q=-\frac{Q}{2}\int\sin(\vartheta)\vartheta^{\prime}\mathop{}\!\mathrm{d}r=-\frac{Q}{2}\int_{\pi}^{0}\sin(\vartheta)\mathop{}\!\mathrm{d}\vartheta=-\frac{Q}{2}[\cos\vartheta]_{0}^{\pi}=Q\ .$ (2.33)
@@
$E=2\pi\int r\mathop{}\!\mathrm{d}r\left(\dfrac{1}{2}(\vartheta^{\prime})^{2}+\dfrac{\sin^{2}\vartheta}{2r^{2}}+\kappa\left(\vartheta^{\prime}+\dfrac{\sin(2\vartheta)}{2r}\right)+h(1-\cos\vartheta)^{p}\right).$ (2.35)
The associated variational equation is
$\vartheta^{\prime\prime}+\dfrac{1}{r}\vartheta^{\prime}+\frac{2\kappa}{r}\sin^{2}\vartheta-\dfrac{1}{2r^{2}}\sin(2\vartheta)-hp\sin\vartheta(1-\cos\vartheta)^{p-1}=0\ .$ (2.36)
@@
2.2 The DM term and integration-by-parts
It is well known that adding a total derivative to the energy does not change the equations of motion, but it may change the energy. The DM term in the canonical formulation with the hedgehog Ansatz (2.31) (with $Q=-1$) is written as
$E_{1}=2\pi\kappa\int r\mathop{}\!\mathrm{d}r\left(\vartheta^{\prime}+\frac{\sin 2\vartheta}{2r}\right),$ (2.39)
but integration by parts as in Eq. (2.16) yields
$E_{1}^{b}=4\pi\kappa\int r\mathop{}\!\mathrm{d}r\,\sin^{2}(\vartheta)\vartheta^{\prime}+\pi\kappa[r\sin 2\vartheta]_{0}^{\infty}\ ,$ (2.40)
where the latter is a total derivative. A different integration by parts can be performed, so as to eliminate $\vartheta^{\prime}$,
$E_{1}^{c}=2\pi\kappa\int r\mathop{}\!\mathrm{d}r\left(-\frac{\vartheta}{r}+\frac{\sin 2\vartheta}{2r}\right)+2\pi\kappa[r\vartheta]_{0}^{\infty}\ .$ (2.41)
If $p=1$ the total derivatives simply vanish and they can safely be dropped without changing the energies. However, for the $p>1$ cases, all three DM terms without the boundary terms give rise to the same equation of motion, but different energies if the boundary terms are dropped.
We will see shortly, that we need to drop a boundary term and compute the DM energy with a modified term in order to predict the correct properties of the physical solution in the $\operatorname{O}(3)$ sigma model limit, which we shall study next.

### v3

@@ Abstract
The magnetic Skyrmion is described by one control parameter and one length scale. We study the two extreme limits of the control parameter – infinitely large and vanishing – and find that the magnetic Skyrmion becomes a “restricted” magnetic Skyrmion and an $\operatorname{O}(3)$ sigma model lump, respectively. Depending on the potential under consideration, the restricted limit manifests differently. In the case of the Zeeman term, the restricted magnetic Skyrmion becomes a “supercompacton” that develops a discontinuity, whereas for the Zeeman term to the power 3/2 it becomes a normal compacton. In both the lump and the restricted limit the solution is given in exact explicit form. We observe that the case of the Zeeman term squared, which can also be understood as a special combination of the Zeeman term and the easy-plane potential – realizable in the laboratory, the analytically exact solution for all values of the coupling – including the Bogomol’nyi-Prasad-Sommerfield (BPS) case – is also of the lump type. Finally, we notice that certain materials (e.g. Fe1-xCoxSi or Mn1-xFexGe) have a rather large control parameter $\epsilon$ of order 100, making the restricted limit a suitable rough approximation.
@@ 1 Introduction
In this work we consider the theory of magnetic Skyrmions in two spatial dimensions (2D) and apply certain techniques used in the past in the context of the baby-Skyrme model [42]. We show that a dimensionless parameter enables us to control the relation between the Dirichlet term, the one with two derivatives, and the rest of the energy functional that contains the potential and the DM term. By changing this parameter we can flow from an almost Bogomol’nyi-Prasad-Sommerfield (BPS) theory, close to the $\operatorname{O}(3)$ sigma model which admits lumps as solutions, to a “restricted” magnetic Skyrmion model with just $E_{1}$ and $E_{0}$. The restricted model is analytically solvable. In the case of the ordinary Zeeman potential, which is quadratic in the perturbative excitation fields near the vacuum, a feature never encountered before in the context of the baby-Skyrme model is present: the solution is not only a compacton111 Compactons are solitonic solutions with compact support: they have a nontrivial profile function for $r<R$ and vanishing energy outside, i.e. for $r>R$. Often the fields are continuous but their derivatives are not at $r=R$. Compactons exist in the baby-Skyrme model [43, 44, 45], the BPS-Skyrme model [46], as Q-balls in the gauged (ungauged) $\mathbb{C}P^{N}$ model [47]([48]) and in K-field theories [49], etc. but it is also discontinuous. We may call it a “supercompacton”. The discontinuity is resolved as soon as we turn on a small coefficient for the Dirichlet term. We also confirm the result of Ref. [20], that the magnetic Skyrmion being a solution to the variational equations of motion only corresponds to a stationary point of the energy functional if a certain boundary term is subtracted off. This boundary term can be written in a form of a vorticity [19, 20]. Indeed, this proves important for determining the correct lump sizes in the limit of the control parameter ($\epsilon$) going to zero.
@@
Finally, we observe that this lump-type solution corresponds to the solution of the critically coupled version of magnetic Skyrmions with the Zeeman term squared, as described by Barton-Singer et al[19]. It is not limited to this scenario: it is also the solution to the full second-order variational equation of motion for all values of the coupling, first observed by Döring and Melcher [50]. We find that for this potential, discarding the boundary term changes the energy functional to become dependent on $\epsilon$, hence revealing a phase transition between the homogeneous phase and the Skyrmionic phase.
@@
On the experimental side, we also confront the dimensionless parameter with the various experimental realizations of the magnetic Skyrmion and find that the (near) “restricted” magnetic Skyrmion – corresponding to the case of a large control parameter – is actually realized in some materials, like Fe1-xCoxSi or Mn1-xFexGe. We also confirm that it is possible to experimentally realize the Zeeman squared term, by using magnets with a certain anisotropy term and adjusting the magnetic field of the Zeeman term, so as to complete the square. At this critically tuned value of the magnetic field, the exact analytic lump-type solution is thus realizable in the laboratory.
@@
This paper is organized as follows. In Sec. 2 we review the magnetic Skyrmion in the continuum formulation with various potentials in our notation. In Sec. 2.1, we review the hedgehog Ansatz and discuss the symmetry breaking properties leading to the fixing of the phase parameter. In Secs. 2.2 and 2.3 we introduce the sigma model lump and restricted limits of the magnetic Skyrmion. In Sec. 2.4 we review the case of the critical coupling and in Sec. 2.5 we show that the lump solution is a solution to the model with the Zeeman term squared for any coupling. In Sec. 2.6 we provide numerical solutions that interpolate between the lump and restricted limits, verifying the analytical results obtained. In Sec. 2.7 we compare the control parameter to experimental parameters of various chiral magnets. We conclude in Sec. 3 with a discussion.
@@
$\mathcal{H}_{\mathrm{H}}=-\sum_{i<j}\mathbb{J}_{ij}\ \bm{S}_{i}\cdot\bm{S}_{j},$ (2.1)
where the interaction is only between nearest neighbor’s spins, and $\mathbb{J}_{ij}$ is the exchange interaction constant (which is negative in ferromagnetic systems). We suppose that $\mathbb{J}_{ij}=J$ for $i$ and $j$ being nearest neighbor sites in the crystal. In the mean field approximation, a proper dynamical variable is the expectation value of the spins, i.e. the unit magnetization vector $\bm{n}\in S^{2}$. Its dynamics is governed by the continuum Ginzburg-Landau (GL) effective theory (see e.g. Ref. [6]). This is the static energy of the $\operatorname{O}(3)$ sigma model. Furthermore, the Hamiltonian of the Heisenberg model can be supplemented by the potential-like term $-\bm{B}\cdot\bm{n}$ that describes the Zeeman interaction with an external magnetic field, which we choose to be aligned along the third Cartesian axis, $\bm{B}=B\bm{e}_{3}$.
@@
$\mathcal{H}_{\mathrm{DM}}=-\sum_{ij}\bm{D}_{ij}\cdot(\bm{S}_{i}\times\bm{S}_{j}),$ (2.2)
where $\bm{D}_{ij}$ is the chiral coupling parameter, chosen here to be equal to $D\hat{\bm{e}}_{j-i}$ for $i$ and $j$ being nearest neighbor sites in the crystal, $\hat{\bm{e}}_{j-i}$ is a unit vector pointing from the lattice site $i$ to the lattice site $j$ and $D$ depends on the material. [^2] The chosen DM term here is denoted the Bloch-type DM term in the literature [6, 19, 21]. The corresponding energy in the continuum limit (see App. A), at the leading order in the long-wavelength approximation, reduces to the functional:
@@
$E[\bm{n}]=\int\mathop{}\!\mathrm{d}^{2}x\bigg{(}$ $\frac{1}{2}\partial_{i}\bm{n}\cdot\partial_{i}\bm{n}+\kappa\epsilon_{iab}\partial^{i}n^{a}n^{b}+V_{p}(\bm{n})$
$+\frac{\lambda}{2}(\bm{n}\cdot\bm{n}-1)\bigg{)},$ (2.4)
@@
As was shown in Ref. [20], the energy functional is only stationary with respect to the solutions to the variational equations of motion if a certain boundary term is subtracted off:
$\widetilde{E}[\bm{n}]$ $=E[\bm{n}]-\Omega[\bm{n}],$ (2.7)
$\Omega[\bm{n}]$ $=\kappa\int\mathop{}\!\mathrm{d}^{2}x\;\epsilon_{iab}\partial^{i}n^{a}N^{b},$ (2.8)
where $\Omega[\bm{n}]$ is called the vorticity. The vorticity is not part of the DM term, but a similar term that is locally different, yet giving the same boundary integral for finite-energy configurations (i.e. configuration with $\lim_{|\bm{x}|\to\infty}\bm{n}=\bm{N}$) is
$\widehat{\Omega}[\bm{n}]=\kappa\int\mathop{}\!\mathrm{d}^{2}x\;\epsilon_{ia3}\partial^{i}(n^{a}n^{3}).$ (2.9)
Subtracting either the vorticity $\Omega$ or the boundary term $\widehat{\Omega}$ from the energy functional $E[\bm{n}]$ will give rise to a well-defined variational problem. The first option is given in Eq. (2.7) as $\widetilde{E}[\bm{n}]$ and we will call the second option $\widehat{E}[\bm{n}]=E[\bm{n}]-\widehat{\Omega}[\bm{n}]$.
@@
So the meaning of $\bm{T}$ is a pair of orthogonal unit vectors in $\mathbb{R}^{3}$. The energy functional (2.3) corresponds to the choice
$\bm{T}=\begin{pmatrix}1&0\\ 0&1\\ 0&0\end{pmatrix},\qquad\bm{N}=\begin{pmatrix}0\\ 0\\ 1\end{pmatrix},$ (2.12)
which in indices is $T^{ai}=\delta^{ai}$ and corresponds to the Bloch-type DM term. A more general DM term can thus be written using a different tensor obeying the conditions (2.11). In particular, Refs. [19, 21] use
$\bm{T}=\begin{pmatrix}\cos\alpha&-\sin\alpha\\ \sin\alpha&\cos\alpha\\ 0&0\end{pmatrix},$ (2.13)
where $\alpha=0$ corresponds to the Bloch-type DM term and $\alpha=\frac{\pi}{2}$ to the Néel-type DM term.
The equations of motion are independent of boundary terms and read
$\partial_{i}^{2}n^{a}+hpN^{a}\left(1-\bm{n}\cdot\bm{N}\right)^{p-1}-2\kappa\epsilon_{aib}\partial^{i}n^{b}-\lambda n^{a}=0,$ (2.14)
$\bm{n}\cdot\bm{n}=1$, where the Lagrange multiplier is
@@
$\widetilde{E}[\vartheta,\varphi]$ $=\int\mathop{}\!\mathrm{d}^{2}x\,\bigg{(}\frac{1}{2}(\partial_{i}\vartheta)^{2}+\frac{1}{2}\sin^{2}\vartheta(\partial_{i}\varphi)^{2}+h(1-\cos{\vartheta})^{p}$
$+\kappa(1-\cos\vartheta)\Big{[}\sin\varphi(\partial_{1}\vartheta-\sin\vartheta\partial_{2}\varphi)$
@@
whereas for reference, the energy functional with the boundary term (2.9) subtracted off instead simplifies as
$\widehat{E}[\vartheta,\varphi]$ $=\int\mathop{}\!\mathrm{d}^{2}x\,\bigg{(}\frac{1}{2}(\partial_{i}\vartheta)^{2}+\frac{1}{2}\sin^{2}\vartheta(\partial_{i}\varphi)^{2}+h(1-\cos{\vartheta})^{p}$
$\quad+\kappa\sin^{2}\vartheta\,\left(\sin\varphi\,\partial_{1}\vartheta-\cos\varphi\,\partial_{2}\vartheta\right)\bigg{)}.$ (2.20)
The simplification occurs because the boundary term (2.9) is already contained in the original DM term of Eq. (2.4).
@@
where $\ell$ controls the size of the Skyrmion (2.22) while $\epsilon$, the dimensionless parameter of the theory, controls the relation between $E_{2}$ and $E_{1}+E_{0}$. We call the latter the control parameter of the theory (i.e. $\epsilon$). In order to see the scaling behavior of the theory, we can write
@@
$\bm{n}(x_{1},x_{2})\mapsto R_{12}(\zeta)\bm{n}(\cos\zeta x_{1}-\sin\zeta x_{2},\sin\zeta x_{1}+\cos\zeta x_{2}),$ (2.29)
$\zeta\in[0,2\pi)$. Moreover, there are the two internal parities as a discrete group $P_{\rm spt}\times P_{\rm int}$. The DM interaction is responsible for the breaking to the diagonal group $P_{\rm diag}$. Up to an additional internal rotation of $\pi$, the parity transformations act as:
@@
$G=\operatorname{O}(2)_{\rm diag}\ltimes T_{2},$ (2.31)
where $T_{2}$ corresponds to the 2-dimensional translations, $\operatorname{O}(2)_{\rm diag}$ contains the diagonal parity and $\ltimes$ denotes a semidirect product.
@@ 2.1 Hedgehog solution
$+\frac{\kappa}{\pi(Q+1)}\sin(Q\pi)\sin(\delta-Q\pi)\left(\vartheta^{\prime}-\dfrac{Q}{2r}\sin(2\vartheta)\right)\bigg{\}},$ (2.33)
@@
$Q$ $=-\frac{Q}{2}\int\sin(\vartheta)\vartheta^{\prime}\mathop{}\!\mathrm{d}r=-\frac{Q}{2}\int_{\pi}^{0}\sin(\vartheta)\mathop{}\!\mathrm{d}\vartheta$
$=-\frac{Q}{2}[\cos\vartheta]_{0}^{\pi}=Q.$ (2.34)
@@
$E=2\pi\int r\mathop{}\!\mathrm{d}r\bigg{(}$ $\dfrac{1}{2}(\vartheta^{\prime})^{2}+\dfrac{\sin^{2}\vartheta}{2r^{2}}+\kappa\left(\vartheta^{\prime}+\dfrac{\sin(2\vartheta)}{2r}\right)$
$+h(1-\cos\vartheta)^{p}\bigg{)},$ (2.36)
Figure 1: The magnetization vector (2.18) for a typical magnetic Skyrmion solution.
and the vector plot of the corresponding magnetization vector (2.18) is shown in Fig. 1. The energy functional with the vorticity and boundary term subtracted off reads
$\widetilde{E}$ $=2\pi\int r\mathop{}\!\mathrm{d}r\bigg{(}\frac{1}{2}(\vartheta^{\prime})^{2}+\dfrac{\sin^{2}\vartheta}{2r^{2}}$
$+\kappa(1-\cos\vartheta)\left(\vartheta^{\prime}-\dfrac{\sin\vartheta}{r}\right)+h(1-\cos\vartheta)^{p}\bigg{)},$ (2.37)
$\widehat{E}$ $=2\pi\int r\mathop{}\!\mathrm{d}r\bigg{(}\frac{1}{2}(\vartheta^{\prime})^{2}+\dfrac{\sin^{2}\vartheta}{2r^{2}}+\kappa\sin^{2}(\vartheta)\vartheta^{\prime}$
$+h(1-\cos\vartheta)^{p}\bigg{)},$ (2.38)
respectively. The associated variational equation is
$\vartheta^{\prime\prime}+\dfrac{1}{r}\vartheta^{\prime}+\frac{2\kappa}{r}\sin^{2}\vartheta-\dfrac{1}{2r^{2}}\sin(2\vartheta)\qquad$
$-hp\sin\vartheta(1-\cos\vartheta)^{p-1}=0.$ (2.39)